OpenStudy (anonymous):

The line with slope 7/9 and containing the midpoint of the segment whose endpoints are (2, -3) and (-6, 5).

OpenStudy (anonymous):

i tryed [y-5=9/7(x+6)]7 7y-35=9x+54 -9X -9x >>>-9x+7y=89

OpenStudy (anonymous):

i also tryed it with the (2,-3) and i got -9x+7y=39

OpenStudy (anonymous):

i keepon getting different ansers but somehow i also keep on getting -9x+7y

OpenStudy (anonymous):

what step am i doing wrong can some1 tell me

OpenStudy (anonymous):

Find the midpoint first.

OpenStudy (anonymous):